# SPECIAL RELATIVITY

### v9

@@ 4.10 Spatiotemporal Transformation of the Transmitted Light Beam
Let us suppose that transmitted light pulse propagates in the $x-z$ plane. Now we are interested in the space-time intensity distribution in this plane. Spatiotemporal coupling arises naturally in transmitted radiation behind the screen, because the transmission process involves the introduction of an angular-frequency dispersion of the transmitted radiation. The emitted light beam is represented with sufficient accuracy as the product of factors separately depending on space and time. However, when the manipulation of the emitted light requires the transmission through a hole in a moving opaque screen, such assumption fails.
@@ 4.11 Applicability of the Ray Optics
Let us move on to consider the predictions of the existing aberration of light theory in the case of a transversely moving screen which has a hole in it (Fig. 7). According to conventional theory, the spot of the transmitted light beam on the observer screen also moves with the same velocity as the moving screen. It is important at this point to emphasize that the electrodynamics dictates that this would also lead to a consequent introduction of an angular-frequency dispersion. However, such angular spectrum change would mean a correction to a deviation of the energy transport direction of the transmitted light beam so that there is a glaring conflict with the prediction of the energy transport direction according to conventional aberration of light theory. The absence of the group velocity along the moving direction is the prediction of conventional aberration of light theory and is obviously absurd from the viewpoint of electrodynamics.
Figure 10: Transverse moving screen which has a hole in it. Light corpuscles are falling normally on the screen and, according to literature, generates an oblique light beam.
@@
This incorrect statement is a straightforward consequence of the generally accepted way of looking at the aberration of light phenomena of most authors of the texbooks. Today one is told that the phenomenon of aberration of light could be interpreted, using corpuscular model of light. Light corpuscles are falling normally on the moving screen and, according to literature, generate oblique light beam as Fig. 10 shows. This wrong argument persists to this day. If the optical system is spatially coherently illuminated, then a satisfactory treatment of the aberration of light should be based on the electromagnetic wave theory.
@@
Some experts believe that the applicability of corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics. Let us see what happens, according to the ray optics, in our case of interest. Light rays are falling normally on the moving screen and generate oblique ray beam as Fig. 10 shows (6). In this situation, we just treat light rays like little particles and the effect is entirely familiar. We would like now to discuss the region of applicability of ray optics. The situation relating to use the ray optics in the theory of aberration of light is complicated. One could naively expect that the region of applicability of ray optics, following from the textbooks reasoning, should be identified with any spatially incoherent radiation. However, incorrect results are obtained by doing so. In particular, a spatially completely incoherent source (e.g. an incandescent lamp or a star) is actually a system of elementary (statistically independent) point sources with different offsets. Radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources. An elementary source produces in front of a hole aperture effectively a plane wave. In other words, the transmission process involves the introduction of an angular frequency dispersion of the transmitted radiation. It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like a deviation of the energy transport direction. This argument gives reason why ray optics is not applicable in the theory of aberration of light from (spatially) completely incoherent sources.
@@
We will illustrate the applicability of ray optics in the theory of aberration of light for a particular class of spatially incoherent light beams. In fact, the ray beam shown in Fig. 10 can be realized as follows. Such beam may be produced by many commonly used lasers with a random spread of phases. One of the very useful properties of laser sources is their ability to produce fields that are highly directional.The intensity of such fields is concentrated in a very narrow solid angle. A method can be proposed for generating ray beam from primary sources by the use of array of randomly phased lasers. Such planar source generates rays which are falling normally on the moving screen (within the laser Rayleigh range) and generate oblique transmitted ray beam, Fig. 10. Intuitively, a transversely moving screen containing a hole acts like a switcher for lasers. Surely, a luminous spot moving at velocity $v$ can be realized simpler, so to speak, ”manually”. We can arrange the laser-like sources along the $x$ axis and switch them on one after another (independently) from left to right with a given time lag. Naturally, we can get a luminous spot moving at any velocity (even at $v>c$). From this example it is seen that in this process no information can be transmitted (along the $x$ axis) since each source radiates independently.
4.12 Measure of Aberration
[…]
@@ 4.13 Moving Large Aperture Emitter
Let us now consider the case when a ”plane-wave” emitter in the lab inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis. An emitter with finite aperture is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. We already know from our discussion from very beginning of this chapter that there is a deviation of the energy transport for the coherent light radiated by the transversely moving emitter, which is nothing else but well-known result: there is the aberration of light from the transversely moving emitter in the inertial frame of reference (Fig. 1).
@@
The specific of our case of interest with the viewpoint of kinematics is that here the screen is at rest in the lab frame and the emitter is moving tangentially to its surface with constant speed with respect to the lab frame. For simplicity we shall assume that the transverse size of the moving ”plane-wave” emitter is very large relative to the transverse size of the hole in the screen. Suppose that an observer, which is at rest with respect to the screen performs the direction of the energy transport measurement, Fig.11.
Figure 11: Aberration of light in an inertial frame of reference. The large aperture ”plane-wave” emitter moving tangentially to its surface. As one of the consequences of the Doppler effect, we find group velocity of the light waves radiated off a large aperture moving emitter. The screen is at rest and we have actually the problem of steady-state transmission. The transmitted light beam is going vertically because it has lost its horizontal (group) velocity component.
@@
If the screen is at rest and the source of light is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In our previous discussion we learned that the screen at rest must be described by Maxwell’s electrodynamics. In other words, the dispersion equation in the Lorentz coordinatization should be applied and kept in a consistent way for both incoming and scattered waves. A dispersion equation in the case of Maxwell’s electrodynamics is $k_{x}^{2}+k^{2}_{y}+k^{2}_{z}-\omega^{2}=0$. The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$ axis the dispersion equation in the absolute time coordinatization will have the same diagonal form $k^{2}_{z}-\omega^{2}=0$ for the (normal) incident wave. At close look at the physics of this subject shows that in the inertial lab frame, where the screen is at rest, we have actually the problem of steady-state transmission. The Doppler effect is absent and the transmitted beam is going vertically because it has lost its horizontal (group velocity) component. That is the transmission appears as shown in Fig.11. We only wish to emphasize here the following point. When the light passes through the small aperture hole we have a light beam whose fields have been perturbed by diffraction, and now not include information about emitter motion.
@@ 4.14 A Point Source in an Inertial Frame of Reference
Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer (far-zone) diffraction. In Fraunhofer diffraction, the phase of the wave is assumed to vary linearly across the detector aperture. This would occur if, for example, a plane wave were incident on the aperture at an angle with respect to the optical axis. In the Fresnel diffraction, we replace the assumption of a linear phase variation with quadratic phase variation. In the far zone, both types of sources produce in front of pupil detection effectively a plane wave. In other words, there is always the physical influence of the instrument on the measurement of the aberration of light in the Fraunhofer zone.
@@
Figure 12: Aberration of light in an inertial frame of reference. A point source produces in front of a hole aperture effectively a plane wave. If the motion of the point source is parallel to the screen, transmitted beam is going vertically. The aberration of light phenomenon is absent in this situation.
@@
One of the important conclusions of the discussion presented above (see Fig. 11 ) is that the aberration of point source is absent in an inertial frame, Fig. 12. There are a number of remarkable effects which are a consequence of the fact that the information about a point source motion is not included into the light beam transmitted through the hole. In fact, this is the key to the binary star paradox discussed in the Chapter 6.
For the rest of our discussion of the aberration of light phenomena in an inertial frame, it will be more convenient if we consider a somewhat modified source of the ”plane wave” emitter type. We have chosen the model described in the Section 4.2 because it is relatively simple. It is sufficiently complicated that it can stand as a prototype which can be generalized for the description of the aberration of light phenomena. No one has ever done all of ”thought experiments” we described in just this way, but we know what would happen from the laws of special relativity which are, of course, based on other experiments.
@@
At this point, a reasonable question arises: why we are not discuss about the aberration of light radiated by a single laser. At first glance, the output light beam shown in Fig. 13 may be produced by a laser. Indeed, one of the specific properties of a laser is ability to produce coherent light beams that are highly directional. We will not be able actually to get into the subject until later. We will analyze the aberration of light transmitted from the laser resonator which is accelerated in an inertial frame in the Section 8.6 and 10.4.
@@ 4.15 The Difference between the Aberration of Light and Sound
We are now finished with the subject of the aberration of light in the initial inertial frame, and will soon go on to discuss the optical phenomena in the accelerated frame. But before doing so, we would like to discuss the optical-acoustic analogies. Sound is a periodic motion of the air caused by vibrations of parts of the source. The explanation of the effect of sound aberration is actually based on the use of the model of single plane-wave emitter. As a model, we use a two-dimensional array of identical coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$. We take the elementary sources to start radiation waves simultaneous.
@@
We shall discuss the situation where there is a finite aperture mirror moving tangentially to its surface. For simplicity, we shall assume that the transverse size of the moving mirror is very small relative to the transverse size of the ”plane-wave” emitter. The observer describes the sound beam from the emitter at rest (in the atmosphere frame) using usual wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the sound speed. This emitter will radiate plane wavefront in a vertical $z$ direction. The wave equation of the moving mirror in the atmosphere frame is the same as wave equation of the emitter at rest. We may consider the amplitude of the beam radiated by plane of oscillating sources as a whole to be resultant of radiated spherical waves. This is because diagonal wave equation has no intrinsic anisotropy. The elementary sources lying on the plane of simultaneity gives rise to spherical radiated wavelets, and these combine according to Huygens principle to form what is effectively a radiated wave. The wavefront emitted by the plane-wave emitter at given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. According to the theory of sound, there is no deviation of the energy transport for the reflected sound beam. A plane wave is falling normally on the finite aperture mirror, and generates a reflected oblique beam. Note that the phase and group velocities are always equal in the atmosphere frame.
@@
According to textbooks, there is no difference between the aberration of light (in the first order in $v/c$) and sound. The geometry of moving mirror reflection is identical, Fig. 2. We now consider the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen. Consequently, it is generally believed that the geometry of the energy transport for light transmitted through a hole in the moving opaque screen is the same, Fig. 7. In the next chapter (see section 5.14) we shall continue our discussion of the aberration of sound theory.
@@
The textbook approach to the aberration of light usually forces the physicist to believe that there are important parallels between the aberration of sound and light in the first order approximation. In fact, the wave equation of the moving mirror (screen) is the same (i.e. diagonal wave equation) as wave equation of the emitter at rest. On the other hand, in the sound theory we talk about Newton space and time rather than about space-time with pseudo-Euclidean geometry. We come to the conclusion that the standard analysis of the aberration of light in the single inertial frame does not take into account the fundamental difference between velocity of light and velocity of sound.
@@
The satisfactory treatment of the relativistic modification should be based on the space-time geometric approach. The usual Minkowski metric is valid in inertial frame where observer is at rest and the procedure of slow clock transport (or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest) is used to assign values to the time coordinate. In this (Lorentz) coordinate system Minkowski metric describes the electrodynamics of the light source at rest with viewpoint of the inertial observer measurements. Now we consider the case when mirror in the inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. The method of synchronization consists in keeping the same set of synchronized clocks used. This synchronization convention preserves the Minkowski metric of the inertial observer and the light source at rest. In this case we are dealing with the Galilean boost of the mirror. We substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric of the inertial observer to obtain Eq. 11. This metric describes the electrodynamics of the moving mirror with viewpoint of the inertial observer measurements. In contrast, the accepted in the previous literature incorrect assumption that an inertial observer (i.e. observer at rest in the inertial frame) and the accelerated ( in the same inertial frame) mirror have common Minkowski metric has always been considered obvious. So that the commonly accepted approach to special relativity in the first order approximation and sound theory are really just variants of the same classical theory.
4.16 Bibliography and Notes
1. For a general discussion of the Fourier transform methods of spatial filtering theory or Abbe diffraction theory we suggest reading the book [22].
2. It is generally believed that there is no aberration for light reflected from mirrors moving transversely. To quote e.g. Sommerfeld [23]: ”Thus, for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved.” Similar statements can also be found in other textbooks. To quote e.g. Ugarov [24]: ” Hence, when the mirror moves parallel to itself the frequency of incident light is equal to that of reflected light, and the angle of incidence is equal to the angle of reflection.”
3. There is another interpretation that we can give to the scattering process. We note the very close correspondence between angular and frequency response of moving diffraction gratings and the Raman scattering [25]. Since the theory of Bragg diffraction applies to light scattering by sound waves in liquids and solids, it is not surprising that we can obtain our results from the quantum theory of light scattering by phonon. This scattering process is also known as Brillouin scattering. The quantum theory leads to equation $\vec{k}_{s}=\vec{k}_{i}\pm\vec{K}_{\perp}$ by the requirement that the momentum is conserved between interacting photon and phonon. The process of Brillouin scattering is a special case of Raman scattering. The momentum balance for scattering involving the emission or absorption of phonon leads to the Bragg condition. The requirement of conservation of energy leads to the equation $\omega_{s}=\omega_{i}\pm\omega_{phon}$ with $\omega_{s}$, $\omega_{i}$, and $\omega_{phon}$ being the radian frequencies, of the scattered photon, the incident photon, and the phonon. The frequency shift that occurs for light scattering from sound waves comes about as a result of the Doppler effect. Quantum mechanically, it is a consequence of the conservation of energy between the participating particles. We come to the conclusion that, according to the hidden choice of absolute time coordinatization in non-relativistic quantum mechanics, the lab observer actually sees the radiation after reflection as a result of a Galilean boost rather than a Lorentz boost.
4. There is a common misconception that the radiation wavefront orientation has objective meaning. To quote Norton [26]: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront. We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation. Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity. If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame then that must be true in all inertial frame.” This incorrect statement is a straightforward consequence of the generally accepted way of looking at special relativity of most physicists. Accepting the postulate on the constancy of the speed of light one also automatically assumes Lorentz coordinates. According to such limiting understanding of the theory of relativity, it is assumed that only Lorentz coordinatization must be used to map the coordinates of events.
5. Note that spatiotemporal coupling is discussed in literature usually in relation with ultrashort laser pulse propagation through a grating monochromator. Ultrashort laser pulses are usually represented as a products of electric field factors separately dependent on space and time. However, when the manipulation of ultrashort laser pulses requires the propagation through a grating monochromator, such assumption fails. In this situation one has to consider in addition (to phase fronts), planes of constant intensity, that is pulse fronts. A pulse-front tilt can be present in the beam due to propagation through an optical setup incorporating dispersive optical element. In the grating monochromator case, the different spectral components of the out-coming pulse travel in different directions. The electric field of a pulse including angular dispersion can be expressed in the Fourier domain $[k_{x},\omega]$ as $\bar{E}(k_{x}-p\omega,\omega)$, while the inverse Fourier transform from the $[k_{x},\omega]$ to the space-time domain $[x,t]$ can be expressed as $E(x,t+px)$, which is the electric field of pulse with a pulse-front tilt. The tilt angle $\theta_{tilt}$ is given by $\tan\theta_{tilt}=cp$. More specifically $p=dk_{x}/d\omega=kd\theta_{D}/d\omega=\lambda/(c\theta_{D}d)$, where $\lambda=2\pi c/\omega$, $\theta_{D}$ is the diffracted angle, and $d$ is the groove spacing. The diffracted angle $\theta_{D}$ is a function of frequency, according to the well-known plane grating equation. Assuming diffraction into the first order, one has $\lambda=(\cos\theta_{i}-\cos\theta_{D})d$, where $\theta_{i}$ is the incident angle. By differentiating this equation one obtains $d\theta_{D}/d\lambda=1/(\theta_{D}d)$, where we assume for simplisity grazing incidence geometry, $\theta_{i}\ll 1$ and $\theta_{D}\ll 1$. The physical meaning of the former equation is that different spectral components of the out-coming pulse travel in different directions. Therefore one concludes that the pulse-front tilt is invariably accompanied by angular dispersion. It follows that any device like a grating monochromator, that produces an angular dispersion, also introduces significant pulse-front tilt. In our case of interest we have deal with light diffracted by a (transversely) moving set of gratings and the Doppler effect is responsible for frequency dispersion $d\omega/dk_{x}=v$. Thus the spatitemporal coupling due to the light beam transmission through a hole in a moving screen and the usual pulse-front tilt distortion are quite different.
6. To quote Brillouin [27]: ” Fig. 5 explain the situation assuming a simplified device consisting of parallel plate moving with uniform velocity $v$ in the horizontal direction. Monochromatic light is falling normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in our Fig. 10.
@@ 4.16 Bibliography and Notes
5.1 Introductory Remarks
In this section we reexamine the issue of light transmission in non-inertial frames with particular reference to the aberration of light phenomena. We derive the aberration for a pulse of light traveling in the accelerated systems using the Langevin metric in general relativity (1). It comes out naturally if one writes the equation of the time transfer, from the inertial frame to the accelerated frame, in a generally covariant context. We only wish to emphasize here the following point. From a mathematical standpoint, there is no difference between calculations in the framework of the general theory of relativity and of the special theory of relativity in the absence of space-time curvature.
The aberration of light problem is solved with discovery of the essential asymmetry between the non-inertial and the inertial observers. Actually, in resent years it seems to be almost normally accepted in scientific community that the ”theory of relativity” is just a name, not to be taken literally. One can conclude that not all is relative in relativity, because this theory also contains some features which are absolute.
Has all our talk about asymmetry violated the relativity principle? At a first glance it might seem so, since the relativity principle is often interpreted as implying perfect symmetry among moving frames. The principle of relativity denied the possibility for an observer partaking in a uniform motion relative to an inertial frame of discovering by any measurement such a motion, of course, that one does not look outside. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and non-inertial reference systems are not equitable.
A typical resolution of the asymmetry paradox identifies acceleration as the agency of asymmetry. It should be note that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame, so the acceleration need not be considered explicitly in working the problem. Nevertheless, the acceleration completely determines the problem. It is worth remarking that the absent of a dynamical explanation for the asymmetry in special relativity has disturbed some physicists. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer. An idea of treating the accelerated frame using the equivalence principle is useful from a pedagogical point of view. See section 9.1 for more ideas along this line.
However, there remains an intriguing puzzle to solve: how can the (e.g. earth-based and sun-based) observers tell which observer took the acceleration? The surprising fact is, the determination of which observer took the acceleration can be made only by observation of the ”fixed stars”. The acceleration is in principle defined in terms of motion relative to the fixed stars, and they must be consulted in order to determine whether an acceleration occurred. Thus when we state that the earth-based observer undergoes an acceleration, and the sun-based observer does not, there is a hidden assumption concerning the distribution of mass in the universe. The implicit ”absolute” acceleration means acceleration relative to the fixed stars.

### v10

@@ 4.9 Spatiotemporal Transformation of the Transmitted Light Beam
Let us suppose that the transmitted light pulse propagates in the $x-z$ plane. Now we are interested in the space-time intensity distribution in this plane. Spatiotemporal coupling arises naturally in transmitted radiation behind the screen because the transmission process involves the introduction of an angular-frequency dispersion of the transmitted radiation. The emitted light beam is represented with sufficient accuracy as the product of factors separately depending on space and time. However, when the manipulation of the emitted light requires the transmission through a hole in a moving opaque screen, such an assumption fails.
@@ 4.10 Applicability of the Ray Optics
Today one is told that the phenomenon of aberration of light could be interpreted, using a corpuscular model of light. Light corpuscles falls normally on the moving screen and, according to literature, generate oblique light beam as Fig. 10 shows. This wrong argument persists to this day. If the optical system is spatially coherently illuminated, then a satisfactory treatment of the aberration of light should be based on the electromagnetic wave theory.
Figure 10: Tangentially moving screen which has a hole in it. Light corpuscles fall normally on the screen and, according to literature, generate an oblique light beam.
@@
Some experts believe that the applicability of the corpuscular model in the theory of light should be reinterpreted as the applicability of ray optics. Let us see what happens, according to the ray optics, in our case of interest. Light rays fall normally on the moving screen and generate oblique ray beam as Fig. 10 shows  [^23] To quote Brillouin [27]: ” Fig. 5 explain the situation assuming a simplified device consisting of the parallel plate moving with uniform velocity $v$ in the horizontal direction. Monochromatic light is falling normally on the plate and generates an oblique ray.” This oblique effect is demonstrated in Fig. 10.. In this situation, we just treat light rays like little particles and the effect is entirely familiar. We would like now to discuss the region of applicability of ray optics. The situation relating to the use of ray optics in the theory of aberration of light is complicated. One could naively expect that the region of applicability of ray optics, following the textbook’s reasoning, should be identified with any spatially incoherent radiation. However, incorrect results are obtained by doing so. In particular, a spatially completely incoherent source (e.g. an incandescent lamp or a star) is a system of elementary (statistically independent) point sources with different offsets. A radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources. An elementary source produces in front of a hole aperture effectively a plane wave. In other words, the transmission process involves the unavoidable diffraction of a plane wave by an aperture. It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like a deviation of the energy transport direction. This argument gives the reason why ray optics is not applicable to the theory of aberration of light from (spatially) completely incoherent sources.
@@
We will illustrate the applicability of ray optics in the theory of aberration of light for a particular class of spatially incoherent light beams. A method can be proposed for generating ray beam from primary sources by the use of an array of randomly phased lasers. Such a planar source generates rays which are falling normally on the moving screen (within the laser Rayleigh range) and generate an oblique transmitted ray beam, Fig. 10. Intuitively, a tangentially moving screen containing a hole acts like a switcher for lasers. Surely, a luminous spot moving at velocity $v$ can be realized more simply, so to speak, ”manually”. We can arrange the laser-like sources along the $x$ axis and switch them one after another (independently) from left to right with a given time lag. Naturally, we can get a luminous spot moving at any velocity (even at $v>c$). From this example, it is seen that in this process no information can be transmitted (along the $x$ axis) since each source radiates independently.
@@ 4.11 Moving Large Aperture Emitter
Let us now consider the case when a ”plane-wave” emitter in the lab inertial frame is accelerated from rest up to velocity $v$ along the $x$ axis. An emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. We already know from our discussion at the very beginning of this section that there is a deviation of the energy transport for the coherent light radiated by the tangentially moving emitter, which is nothing else but a well-known result: there is the aberration of light from the tangentially moving emitter in the inertial frame of reference (Fig. 1).
@@
The specific of our case of interest from the viewpoint of kinematics is that here the screen is at rest in the lab frame and the emitter is moving tangentially to its surface with constant speed with respect to the lab frame. For simplicity, we shall assume that the size of the moving ”plane-wave” emitter is very large relative to the size of the hole in the screen. Suppose that an observer, which is at rest with respect to the screen performs the direction of the energy transport measurement, Fig.11.
Figure 11: Aberration of light in an inertial frame of reference. The large aperture ”plane-wave” emitter moves tangentially to its surface. The screen is at rest and the aberration of light phenomenon is absent in this situation.
@@
If the screen is at rest and the source of light is in motion, it is obvious that the electrodynamics equations must be identical for all electromagnetic waves. In our previous discussion, we learned that the screen at rest must be described by Maxwell’s electrodynamics. In other words, the dispersion equation in the Lorentz coordinatization should be applied and kept in a consistent way for both incoming and scattered waves. A dispersion equation in the case of Maxwell’s electrodynamics is $k_{x}^{2}+k^{2}_{y}+k^{2}_{z}-\omega^{2}=0$. The peculiarity of the discussed geometry is that even after the Galilean transformation along the $x$ axis the dispersion equation in the absolute time coordinatization will have the same diagonal form $k^{2}_{z}-\omega^{2}=0$ for the (normal) incident wave. A close look at the physics of this subject shows that in the inertial lab frame, where the screen is at rest, we have the problem of steady-state transmission. The Doppler effect is absent and the transmitted beam is going vertically because it has lost its horizontal (group velocity) component. That is the transmission appears as shown in Fig.11. We only wish to emphasize here the following point. When the light passes through the small aperture hole we have a light beam whose fields have been perturbed by diffraction, and now do not include information about emitter motion.
@@ 4.12 A Point Source in an Inertial Frame of Reference
Source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and the Fraunhofer (far-zone) diffraction. In Fraunhofer diffraction, the phase of the wave is assumed to vary linearly across the detector aperture. This would occur if, for example, a plane wave were incident on the aperture at an angle with respect to the optical axis. In the Fresnel diffraction, we replace the assumption of a linear phase variation with a quadratic phase variation. In the far zone, both types of sources produce in front of pupil detection effectively a plane wave. In other words, there is always the physical influence of the instrument on the measurement of the aberration of light in the Fraunhofer zone.
@@
Figure 12: Aberration of light in an inertial frame of reference. In Fraunhofer diffraction region, a point source produces in front of a hole aperture effectively a plane wave. The aberration of light phenomenon is absent in this situation.
@@
One of the important conclusions of the discussion presented above (see Fig. 11 ) is that the aberration of a point source is absent in an inertial frame, Fig. 12. There are a number of remarkable effects that are a consequence of the fact that the information about a point source motion is not included in the light beam transmitted through the hole. In fact, this is the key to the binary star paradox discussed in Chapter 6.
For the rest of our discussion of the aberration of light phenomena in an inertial frame, it will be more convenient if we consider a somewhat modified source of the ”plane wave” emitter type. We have chosen the model described in Section 4.2 because it is relatively simple. It is sufficiently complicated that it can stand as a prototype that can be generalized for the description of the aberration of light phenomena. No one has ever done all of the ”thought experiments” we described in just this way, but we know what would happen from the laws of special relativity which are, of course, based on other experiments.
@@ 4.13 The Difference between the Aberration of Light and Sound
We are now finished with the subject of the aberration of light in the initial inertial frame, and will soon go on to discuss the optical phenomena in the accelerated frame. But before doing so, we would like to discuss the optical-acoustic analogies. Sound is a periodic motion of the air caused by vibrations of parts of the source. The explanation of the effect of sound aberration is based on the use of the model of a single plane-wave emitter. As a model, we use a two-dimensional array of identical coherent elementary sources uniformly distributed on a given $(x-y)$ plane $P$. We take the elementary sources to start radiation waves simultaneously.
@@
We shall discuss the situation where there is a finite aperture mirror moving tangentially to its surface. For simplicity, we shall assume that the size of the moving mirror is very small relative to the size of the ”plane-wave” emitter. The observer describes the sound beam from the emitter at rest (in the atmosphere frame) using the diagonal wave equation $\nabla^{2}f-\partial^{2}f/\partial(v_{s}t)^{2}=0$, where $v_{s}$ is the sound speed. This emitter will radiate plane wavefront in a vertical $z$ direction. The wave equation of the moving mirror in the atmosphere frame is the same as that of the emitter at rest. We may consider the amplitude of the beam radiated by the plane of oscillating sources as a whole to be the result of radiated spherical waves. This is because the diagonal wave equation has no intrinsic anisotropy. The elementary sources lying on the plane of simultaneity give rise to spherical radiated wavelets, and these combine according to Huygens’s principle to form what is effectively a radiated wave. The wavefront emitted by the plane-wave emitter at a given instant time is constructed as an envelope of all the spherical wavelets from the point sources on the emitter that have radiated until this instant. According to the theory of sound, there is no deviation of the energy transport for the reflected sound beam. A plane wave falls normally on the finite aperture mirror and generates a reflected oblique beam. Note that the phase and group velocities are always equal in the atmosphere frame.
@@
According to textbooks, there is no difference between the aberration of light (in the first order in $v/c$) and sound. The geometry of moving mirror reflection is identical, Fig. 2. We now consider the case of a screen, in the atmosphere frame, moving with velocity $v$ along its surface. According to the theory of sound, there is no deviation of the energy transport for sound transmitted through a hole in the moving screen. Consequently, it is generally believed that the geometry of the energy transport for light transmitted through a hole in the moving opaque screen is the same, Fig. 7. In the next section, we shall continue our discussion of the aberration of sound theory.
@@
The textbook approach to the aberration of light usually forces the physicist to believe that there are important parallels between the aberration of sound and light in the first-order approximation. The wave equation of the moving mirror (screen) is the same (i.e. diagonal wave equation) as the wave equation of the emitter at rest. On the other hand, in the sound theory, we talk about Newton’s space and time rather than about space-time with pseudo-Euclidean geometry. We conclude that the standard analysis of the aberration of light in the single inertial frame does not take into account the fundamental difference between the velocity of light and the velocity of sound.
@@
The satisfactory treatment of the relativistic modification should be based on the space-time geometric approach. The usual Minkowski metric is valid in the inertial frame where the observer is at rest and the procedure of slow clock transport (or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest) is used to assign values to the time coordinate. In this (Lorentz) coordinate system Minkowski metric describes the electrodynamics of the light source at rest with the viewpoint of the inertial observer measurements. Now we consider the case when the mirror in the inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. The method of synchronization consists in keeping the same set of synchronized clocks used. This synchronization convention preserves the Minkowski metric of the inertial observer and the light source at rest. In this case, we are dealing with the Galilean boost of the mirror. We substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric of the inertial observer to obtain Eq. 11. This metric describes the electrodynamics of the moving mirror from the viewpoint of the inertial observer measurements. In contrast, the accepted in the previous literature incorrect assumption that an inertial observer (i.e. observer at rest in the inertial frame) and the accelerated ( in the same inertial frame) mirror have a common Minkowski metric has always been considered obvious. So the commonly accepted approach to special relativity in the first-order approximation and sound theory are just variants of the same classical theory.
@@
In this section, we reexamine the issue of light transmission in non-inertial frames with particular reference to the aberration of light phenomena. We derive the aberration for a pulse of light traveling in the accelerated systems using the Langevin metric in general relativity  [^24] The aberration of light problem in the accelerated systems is solved with the discovery of the essential asymmetry between the non-inertial and the inertial observers. This asymmetry is of the same nature as that of the well-known Sagnac effect [28, 29, 30]. For instance, Langevin’s 1921 explanation of the Sagnac effect rested upon the assertion that ”any change of velocity, or any acceleration has an absolute meaning.” [31].. We only wish to emphasize here the following point. From a mathematical standpoint, there is no difference between calculations in the framework of the general theory of relativity and of the special theory of relativity in the absence of space-time curvature.
The aberration of light problem is solved with the discovery of the essential asymmetry between the non-inertial and the inertial observers. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and non-inertial reference systems are not equitable. A typical resolution of the asymmetry paradox identifies acceleration as the agency of asymmetry. It should be noted that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame, so the acceleration need not be considered explicitly in working the problem. Nevertheless, the acceleration completely determines the problem. It is worth remarking that the absence of a dynamical explanation for the asymmetry in special relativity has disturbed some physicists. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer. The idea of treating the accelerated frame using the equivalence principle is useful from a pedagogical point of view. We will discuss this subject further in Chapter 12.
At this point, a reasonable question arises: how can the (e.g. earth-based and sun-based) observers tell which observer took the acceleration? The surprising fact is, the determination of which observer took the acceleration can be made only by observation of the ”fixed stars”. The acceleration is in principle defined in terms of motion relative to the fixed stars. The implicit ”absolute” acceleration means acceleration relative to the fixed stars.